# From Reasoning LLMs to BERT: A Two-Stage Distillation Framework for Search Relevance

### v2

@@ Abstract.
Query-service relevance prediction in e-commerce search systems faces strict latency requirements that prevent the direct application of Large Language Models (LLMs). To bridge this gap, we propose a two-stage reasoning distillation framework to transfer reasoning capabilities from a powerful teacher LLM to a lightweight, deployment-friendly student model. In the first stage, we address the limitations of general-purpose LLMs by constructing a domain-adapted teacher model. This is achieved through a three-step process: domain-adaptive pre-training to inject platform knowledge, supervised fine-tuning to elicit reasoning skills, and preference optimization with a multi-dimensional reward model to ensure the generation of reliable and preference-aligned reasoning paths. This teacher can then automatically annotate massive query-service pairs from search logs with both relevance labels and reasoning chains. In the second stage, to address the challenges of architectural heterogeneity in standard distillation, we introduce Contrastive Reasoning Self-Distillation (CRSD). By modeling the behavior of the same student model under ”standard” and ”reasoning-augmented” inputs as a teacher-student relationship, CRSD enables the lightweight model to internalize the teacher’s complex decision-making mechanisms without needing the explicit reasoning path at inference. Offline evaluations and online A/B testing in the Meituan search advertising system demonstrate that our framework achieves significant improvements across multiple metrics, validating its effectiveness and practical value.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Retrieval models and rankingccs: Information systems Language models
@@ 1. Introduction
In the realm of e-commerce, the core mission of a search engine is to precisely connect user queries with a vast repository of products, making query-service relevance prediction a cornerstone technology for achieving this goal. In large-scale e-commerce search systems, the search process typically employs a multi-stage architecture (Guo et al., 2022; Momma et al., 2022): A retrieval module (coarse ranking) selects a candidate set from the massive item repository; subsequently, a relevance model within the fine-grained ranking stage conducts a detailed evaluation of each ¡query, service¿ pair, filtering out irrelevant results to ensure the quality and user experience of the final-displayed products. As the relevance model must handle hundreds of millions of query-service evaluations daily, it must possess exceptional inference efficiency while maintaining high accuracy(Khattab and Zaharia, 2020).
@@
In recent years, the remarkable capabilities of Large Language Models (LLMs) have offered a promising new direction for tackling these challenges (Wei et al., 2022; Guo et al., 2025; Zhao et al., 2023). Through large-scale pre-training, these models internalize extensive world knowledge, which serves as the foundation for their advanced abilities in natural language understanding and logical reasoning. This foundation enables them not only predict relevance labels, facilitating the large-scale generation of training data to substitute for manual annotation, but can also provide detailed reasoning paths to explain their judgments upon request. This ”reasoning process” or ”chain-of-thought” represents the cognitive steps the model takes to reach a specific decision. Such an ability provides interpretability for relevance prediction and has the potential to yield more accurate and trustworthy outcomes.
@@
Furthermore, while lightweight models like BERT can satisfy latency requirements and are widely used in industry, their heavy reliance on data quality and inherent deficiencies in general knowledge and reasoning capabilities are significant limitations. The prevailing approach to bridge this gap is knowledge distillation—using a teacher model to transfer its knowledge to a small model. However, traditional distillation methods face two key technical obstacles when transferring reasoning abilities: (1) Architectural Heterogeneity: The disparity between decoder-only LLMs and encoder-only BERT models makes direct feature alignment difficult (Hofstätter et al., 2020). (2) Focus on Output Imitation over Process Understanding: Existing methods primarily concentrate on mimicking the final output, failing to effectively transfer the intrinsic logical mechanisms of the reasoning process.
@@
Stage One: Building a Domain-Adapted Reasoning LLM. To address the shortcomings of general-purpose LLMs in platform-specific scenarios, we construct a specialized domain reasoning LLM. This is achieved through a three-step training pipeline: continuous pre-training to inject domain knowledge, supervised fine-tuning to elicit reasoning abilities, and preference optimization using a reward model for fine-grained enhancement of reasoning quality. This model not only overcomes the limitations of generic models but is also capable of generating high-quality relevance labels and detailed reasoning paths for massive volumes of query-service pairs.
@@
We conducted rigorous evaluations on a large-scale, real-world dataset and validated the method’s effectiveness through online A/B testing within Meituan’s e-commerce search system. The results demonstrate that our approach significantly improves relevance prediction accuracy while maintaining inference efficiency, offering a practical solution for reasoning capability transfer in industrial applications.
Our main contributions are threefold:
@@
  • Comprehensive Validation and Production Deployment: We demonstrate the effectiveness of our approach through extensive experiments on real-world datasets and a successful online deployment in a large-scale e-commerce search engine, confirming significant accuracy improvements.
@@ 2.1. Search Relevance in E-commerce
2.2. Large Reasoning Models for Search Relevance
@@
Building upon these reasoning capabilities, several industrial systems have begun integrating LLMs for search relevance prediction. Wang et al. (Wang et al., 2024) show improvements in user preference modeling at Pinterest. However, the high latency and cost of LLM inference limit their real-time deployment in production systems.
@@ 2.3. Knowledge Distillation for Search Relevance
In information retrieval and e-commerce search, knowledge distillation (KD) has been widely adopted to compress heavy cross-encoders into lightweight bi-encoder or dual-tower models (Hofstätter et al., 2020; Yao et al., 2022; Liu et al., 2022). Building on this foundation, several large-scale e-commerce platforms have integrated KD frameworks into production search systems. For instance, ReprBERT (Yao et al., 2022) was successfully applied to Alibaba’s product retrieval pipeline, and Liu et al. (Liu et al., 2022) designed a contextual distillation model to enhance relevance matching in e-commerce product search. Shang et al. (Shang et al., 2025) leveraged LLM-based distillation to enhance Walmart’s e-commerce search, demonstrating that student models distilled from reasoning-rich teachers can achieve strong performance under strict latency constraints. Extending this direction, Ye et al. (Ye et al., 2024) proposed DisRanker, which integrates point-wise and margin-MSE objectives for web-scale ranking distillation. Zhao et al. (Zhao et al., 2025) introduced a multi-dimensional framework that jointly transfers score distributions and chain-of-thought representations to improve interpretability and relevance consistency. Closely related to our work, Agrawal et al. (Agrawal et al., 2025) encoded reasoning paths generated by an external LLM using a separate sentence encoder and aligned them with BERT’s [CLS] vector. However, this approach overlooks that the reasoning LLM may not align well with domain-specific data, leading to suboptimal label accuracy.
@@ 3.1. Framework Overview
We proposed Reasoning-then-Distilling (RTD) framework in this paper, which operates in two primary stages. The first stage is dedicated to building a powerful teacher model $M_{t}$. As can be seen in Figure 1, this is achieved through domain-adaptive continued pre-training, supervised fine-tuning, and preference optimization guided by a multi-dimensional reward model, collectively ensuring that the teacher model delivers not only accurate relevance labels but also trustworthy reasoning chain. Subsequently, the second stage involves training a lightweight student model $M_{s}$ via a contrastive reasoning self-distillation approach, enabling it to internalize the teacher’s complex decision-making mechanisms without explicitly generating intermediate reasoning steps.
@@ 3.2.1. Domain-Adaptive Continued Pre-Training (CPT)
To enhance an open-source LLM with domain-specific knowledge (e.g., platform-specific service information), we performed domain-adaptive continued pre-training. For this purpose, we first constructed a large-scale, multi-task search relevance corpus. The construction of this corpus involved sourcing data from online search logs and then employing several powerful LLMs for automatic annotation. The resulting corpus comprised millions of samples across seven task types, as detailed in  Table 1. Subsequently, we performed continued pre-training on an open-source base LLM using this multi-task corpus, which yielded the model $M_{cpt}$. This model is now equipped with the foundational domain knowledge for complex reasoning tasks.
@@ 3.2.2. Supervised Fine-Tuning (SFT)
[…]
@@
where $N$ is the total number of training samples. This alignment process compels the model to internalize the human-annotated reasoning patterns. Consequently, this model, denoted as $M_{SFT}$, gains the preliminary ability to produce a logical and business-rule-compliant reasoning chain before delivering its final answer.
@@ 3.2.3. Preference Optimization with a Multi-Dimensional Reward Model
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}(\hat{q},a)\sim\mathcal{D}\\ {o_{i}}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)\end{subarray}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\bigg($
$\qquad\min\Big(\rho_{i,t}(\theta)\hat{A}_{i,t},\;\text{clip}(\rho_{i,t}(\theta),{1-\varepsilon},{1+\varepsilon)}\hat{A}_{i,t}\Big)$
(3) $\qquad-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\bigg)\Bigg]$
@@
(4) $\rho_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid\hat{q},o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid\hat{q},o_{i,<t})}$
Here, the advantage function $\hat{A}_{i,t}=\frac{R_{\text{final}}^{i}-\text{mean}(\{R_{\text{final}}^{i}\}_{i=1}^{G})}{\text{std}(\{R_{\text{final}}^{i}\}_{i=1}^{G})}$, measures the relative quality of the $i$-th output within its group $G$.
@@ 3.3.1. Baseline Reasoning Distillation Method
Reasoning The user’s query ”Freshly-made cocktail” implies an intent for on-site preparation. The service, a ’Homebar’, not only explicitly mentions ’cocktail set’ but also belongs to the ’Bar’ category. The key inference is that cocktails in a bar context are inherently made-to-order, thus satisfying the ’freshly-made’ requirement… Therefore, the service is highly relevant.
@@ 3.3.2. Contrastive Reasoning Self-Distillation
Based on this, we propose a core hypothesis: if a model learns to process a reasoning-augmented input <query, service, reason> during training, this acquired reasoning pattern can be internalized into the model’s parameters through a suitable mechanism. Ultimately, even when given only the standard input <query, service>, the model can spontaneously apply these logics, achieving a transition from ”explicit reliance on reasoning” to ”implicit autonomous reasoning.”
@@
  • Teacher Configuration: The model receives the reasoning-augmented input <q, s, r>, and its output [CLS] representation is denoted as $\text{cls}_{r}$.
  • Student Configuration: The model receives the standard input <q, s>, and its output [CLS] representation is denoted as cls.
@@
First, we compute standard classification losses for both configurations to ensure the model learns to classify relevance correctly in both modes. We assign a smaller weight $\beta$ to the teacher’s loss to prevent the model from becoming overly reliant on the lengthy reasoning path.
@@
Figure 3. Left: Overview of the proposed CRSD framework, showing the self-distillation process between reasoning-augmented and standard inputs. Right: Illustration of the $\mathcal{L}_{\text{align}}$ objective, which enforces semantic consistency through in-batch contrastive alignment.
Second, and most critically, we introduce an in-batch contrastive learning loss to align the semantic representations from the two configurations. Within a single batch, for the $i$-th sample, its student representation $\text{cls}_{i}$ and teacher representation $\text{cls}_{r_{i}}$ form a positive pair. The student representation $\text{cls}_{i}$ and all other teacher representations $\text{cls}_{r_{j}}$ ($j\neq i$) form negative pairs. By pulling positive pairs closer and pushing negative pairs apart, we compel the model, even with only standard input, to produce a semantic representation similar to the one it would generate if it ”saw” the reasoning path. This loss (InfoNCE) is defined as:
@@
The choice of small values for these hyperparameters is crucial. An overly large $\gamma$ would cause the model to heavily rely on the explicit reasoning path available only in the teacher configuration, failing to learn how to infer relevance from the standard input alone. Similarly, an excessive $\delta$ would prioritize the representation alignment task at the expense of the primary classification objective ($\mathcal{L}_{\text{sce}}$), which is the ultimate goal for the deployed student model. Through this strategy, reasoning ability is effectively ”distilled” from the teacher configuration to the student configuration and ultimately internalized as part of the model itself.

### v3

@@ Abstract.
Query-service relevance prediction in e-commerce search systems faces strict latency requirements that prevent the direct application of Large Language Models (LLMs). To bridge this gap, we propose a two-stage reasoning distillation framework to transfer reasoning capabilities from a powerful teacher LLM to a lightweight, deployment-friendly student model. In the first stage, we address the limitations of general-purpose LLMs by constructing a domain-adapted teacher model. This is achieved through a three-step process: domain-adaptive pre-training to inject platform knowledge, supervised fine-tuning to elicit reasoning skills, and preference optimization with a multi-dimensional reward model to ensure the generation of reliable and preference-aligned reasoning paths. This teacher can then automatically annotate massive query-service pairs from search logs with both relevance labels and reasoning chains. In the second stage, to address the challenges of architectural heterogeneity in standard distillation, we introduce Contrastive Reasoning Self-Distillation (CRSD). By modeling the behavior of the same student model under “standard” and “reasoning-augmented” inputs as a teacher-student relationship, CRSD enables the lightweight model to internalize the teacher’s complex decision-making mechanisms without needing the explicit reasoning path at inference. Offline evaluations and online A/B testing in the Meituan search advertising system demonstrate that our framework achieves significant improvements across multiple metrics, validating its effectiveness and practical value.
@@
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792801ccs: Information systems Retrieval models and rankingccs: Information systems Language models
@@ 1. Introduction
In the realm of e-commerce, the core mission of a search engine is to precisely connect user queries with a vast repository of products, making query-service relevance prediction a cornerstone technology for achieving this goal. In large-scale e-commerce search systems, the search process typically employs a multi-stage architecture (Guo et al., 2022; Momma et al., 2022): A retrieval module (coarse ranking) selects a candidate set from the massive item repository; subsequently, a relevance model within the fine-grained ranking stage conducts a detailed evaluation of each ¡query, service¿ pair, filtering out irrelevant results to ensure quality and user experience of the final-displayed products. As the relevance model must handle hundreds of millions of query-service evaluations daily, it must possess exceptional inference efficiency while maintaining high accuracy (Khattab and Zaharia, 2020).
@@
In recent years, the remarkable capabilities of Large Language Models (LLMs) have offered a promising new direction for tackling these challenges (Wei et al., 2022; Guo et al., 2025; Zhao et al., 2023). Through large-scale pre-training, these models internalize extensive world knowledge, serving as the foundation for their advanced abilities in natural language understanding and logical reasoning (Xu et al., 2025). This foundation enables them not only predict relevance labels, facilitating the large-scale generation of training data to substitute for manual annotation, but also provide detailed reasoning paths to explain their judgments upon request. This “reasoning process” or “chain-of-thought” represents the cognitive steps the model takes to reach a specific decision. Such a reasoning ability provides interpretability for relevance prediction and has the potential to yield more accurate and trustworthy outcomes.
@@
Furthermore, while lightweight models like BERT can satisfy latency requirements and are widely used in industry, their heavy reliance on data quality and inherent deficiencies in general knowledge and reasoning capabilities are significant limitations. The prevailing approach to bridge this gap is knowledge distillation— using a teacher model to transfer its knowledge to a small model. However, traditional distillation methods face two key technical obstacles when transferring reasoning abilities: (1) Architectural Heterogeneity: The disparity between decoder-only LLMs and encoder-only BERT models makes direct feature alignment difficult (Hofstätter et al., 2020). (2) Focus on Output Imitation over Process Understanding: Existing methods primarily focus on mimicking the final output, failing to effectively transfer the intrinsic logical mechanisms of the reasoning process.
@@
Stage One: Building a Domain-Adapted Reasoning LLM. To address the shortcomings of general-purpose LLMs in platform-specific scenarios, we construct a specialized domain reasoning LLM. This is achieved through a three-step training pipeline: continuous pre-training to inject domain knowledge, supervised fine-tuning to elicit reasoning abilities, and preference optimization using a reward model for fine-grained enhancement of reasoning quality. This model not only overcomes the limitations of generic models but also generates high-quality relevance labels and detailed reasoning paths for massive query-service pairs.
@@
We conducted rigorous evaluations on a large-scale, real-world dataset and validated the method’s effectiveness through online A/B testing within Meituan’s e-commerce search system. The results demonstrate that our approach significantly improves relevance prediction accuracy while maintaining inference efficiency, offering a practical solution for reasoning capability transfer in industrial applications. Our main contributions are threefold:
@@
  • Comprehensive Validation and Production Deployment: We demonstrate the effectiveness of our approach through extensive experiments on real-world datasets and a successful online deployment in a large-scale e-commerce search engine.
@@ 2.1. Search Relevance in E-commerce
2.2. Reasoning Models for Search Relevance
@@
Building upon these reasoning capabilitaies, several industrial systems have begun integrating LLMs for search relevance prediction. Wang et al. (Wang et al., 2024) show improvements in user preference modeling at Pinterest. However, the high latency and cost of LLM inference limit their real-time deployment in production systems.
@@ 2.3. Knowledge Distillation for Search Relevance
In information retrieval and e-commerce search, knowledge distillation has been widely adopted to compress heavy cross-encoders into lightweight bi-encoder or dual-tower models (Hofstätter et al., 2020; Yao et al., 2022; Zhang et al., 2022; Liu et al., 2022). Building on this foundation, several large-scale e-commerce platforms have integrated KD frameworks into production search systems. For instance, ReprBERT (Yao et al., 2022) was successfully applied to Alibaba’s product retrieval pipeline, and Liu et al. (Liu et al., 2022) designed a contextual distillation model to enhance relevance matching in e-commerce product search.
Shang et al. (Shang et al., 2025) leveraged LLM-based distillation to enhance Walmart’s e-commerce search, demonstrating that student models distilled from reasoning-rich teachers can achieve strong performance under strict latency constraints. Extending this direction, Ye et al. (Ye et al., 2024) proposed DisRanker, which integrates point-wise and margin-MSE objectives for web-scale ranking distillation. Zhao et al. (Zhao et al., 2025) introduced a multi-dimensional framework that jointly transfers score distributions and chain-of-thought representations to improve interpretability and relevance consistency. Closely related to our work, Agrawal et al. (Agrawal et al., 2025) encoded reasoning paths generated by an external LLM using a separate sentence encoder and aligned them with BERT’s [CLS] vector. However, this approach overlooks that the reasoning LLM may not align well with domain-specific data, leading to suboptimal label accuracy.
@@ 3.1. Framework Overview
We proposed Reasoning-then-Distilling framework in this paper, which operates in two primary stages. The first stage is dedicated to building a powerful teacher model $M_{t}$. As can be seen in Figure 1, this is achieved through domain-adaptive continued pre-training, supervised fine-tuning, and preference optimization guided by a multi-dimensional reward model, collectively ensuring that the teacher model delivers not only accurate relevance labels but also trustworthy reasoning chain. Subsequently, the second stage involves training a lightweight student model $M_{s}$ via a contrastive reasoning self-distillation approach, enabling it to internalize the teacher’s complex decision-making mechanisms without explicitly generating intermediate reasoning steps.
@@ 3.2.1. Domain-Adaptive Continued Pre-Training (CPT)
To enhance an open-source LLM with domain-specific knowledge (e.g., platform-specific service information), we perform domain-adaptive continued pre-training. For this purpose, we first construct a large-scale, multi-task search relevance corpus. The construction of this corpus involves sourcing data from online search logs and then employing several powerful LLMs for automatic annotation. The resulting corpus comprises millions of samples across seven task types, as detailed in  Table 1. Subsequently, we performed continued pre-training on an open-source base LLM using this multi-task corpus, which yields the model $M_{cpt}$. This model is now equipped with the foundational domain knowledge for complex reasoning tasks.
@@ 3.2.2. Supervised Fine-Tuning (SFT)
The goal of this stage is to align the base model with human-like reasoning. To this end, we first curated a high-quality dataset of tens of thousands of samples, each comprising a ¡query, service¿ pair and a meticulously human-annotated reasoning chain leading to a final answer.
We then fine-tuned the base model $M_{cpt}$ on this dataset, enabling it to predict the corresponding reasoning chain $r_{i}$ and answer $a_{i}$ jointly from the input pair $(q_{i},s_{i})$, where $q_{i}$ and $s_{i}$ represent the $i$-th query and service, respectively. The model is optimized by minimizing the standard cross-entropy loss over the sequence:
@@
where $N$ is the total number of training samples. For the remainder of this paper, we consistently use $q$ and $s$ to denote query and service.
This alignment process compels the model to internalize the human-annotated reasoning patterns. Consequently, this model, denoted as $M_{SFT}$, gains the preliminary ability to produce a logical and business-rule-compliant reasoning chain before delivering its final answer.
@@ 3.2.3. Preference Optimization with a Multi-Dimensional Reward Model
While supervised fine-tuning instills human-like reasoning patterns, it does not guarantee logical quality or reliability. To address this limitation, we employ reinforcement learning (RL) to further optimize the model’s reasoning capabilities. However, a critical challenge in applying RL is the design of an effective reward function (Ng et al., 1999; Randløv and Alstrøm, 1998; Tu et al., 2025). A naive approach, such as rewarding only the final answer’s correctness, suffers from two critical issues. Firstly, the reward signal is exceedingly sparse, providing insufficient guidance for the multi-step reasoning process (Zhang et al., 2025). Secondly, it is highly susceptible to “reward hacking,” where the model might generate correct answers through entirely flawed or coincidental reasoning chains, thus failing to achieve the true goal of reliable inference (Tarek and Beheshti, 2025).
@@
(3) $\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}(q,a)\sim\mathcal{D}\\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)\end{subarray}}\Bigg[$ $\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\bigg($
$\min\Big(\rho_{i,t}(\theta)\hat{A}_{i,t},$
$\hskip 18.49988pt\text{clip}\big(\rho_{i,t}(\theta),1-\varepsilon,1+\varepsilon\big)\hat{A}_{i,t}\Big)$
$-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\bigg)\Bigg],$
@@
(4) $\rho_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$
Here, the advantage function
$\hat{A}_{i,t}=\frac{R_{\text{final}}^{i}-\text{mean}(\{R_{\text{final}}^{i}\}_{i=1}^{G})}{\text{std}(\{R_{\text{final}}^{i}\}_{i=1}^{G})}$ measures the relative quality of the $i$-th output within its group $G$.
@@ 3.3.1. Baseline Reasoning Distillation Method
Reasoning The user’s query “Freshly-made cocktail” implies an intent for on-site preparation. The service, a ’Homebar’, not only explicitly mentions ’cocktail set’ but also belongs to the ’Bar’ category. The key inference is that cocktails in a bar context are inherently made-to-order, thus satisfying the ’freshly-made’ requirement… Therefore, the service is highly relevant.
@@ 3.3.2. Contrastive Reasoning Self-Distillation
Based on this, we propose a core hypothesis: if a model learns to process a reasoning-augmented input <query, service, reason> during training, this acquired reasoning pattern can be internalized into the model’s parameters through a suitable mechanism. Ultimately, even when given only the standard input <query, service>, the model can spontaneously apply these logics, achieving a transition from “explicit reliance on reasoning” to “implicit autonomous reasoning.”
@@
  • Teacher Configuration: The model receives the reasoning-augmented input <query, service, reason>, and its output [CLS] representation is denoted as $\text{cls}_{r}$.
  • Student Configuration: The model receives the standard input <query, service>, and its output [CLS] representation is denoted as cls.
@@
First, we compute standard classification losses for both configurations to ensure the model learns to classify relevance correctly in both modes. We assign a smaller weight $\gamma$ to the teacher’s loss to prevent the model from becoming overly reliant on the lengthy reasoning path.
@@
Figure 3. Left: Overview of the proposed CRSD framework. Right: Illustration of the $\mathcal{L}_{\text{align}}$ objective.
Second, and most critically, we introduce an in-batch contrastive learning loss to align the semantic representations from the two configurations. Within a single batch, for the $i$-th sample, its student representation $\text{cls}_{i}$ and teacher representation $\text{cls}_{r_{i}}$ form a positive pair. The student representation $\text{cls}_{i}$ and all other teacher representations $\text{cls}_{r_{j}}$ ($j\neq i$) form negative pairs. By pulling positive pairs closer and pushing negative pairs apart, we compel the model, even with only standard input, to produce a semantic representation similar to the one it would generate if it “saw” the reasoning path. This loss (InfoNCE) is defined as:
@@
We set $\gamma$ and $\delta$ to balance the auxiliary reasoning and alignment tasks with the primary classification goal. This setup ensures the model does not become overly dependent on explicit reasoning paths during training or prioritize representation alignment at the expense of classification accuracy ($\mathcal{L}_{\text{sce}}$). Through this strategy, the reasoning capability is effectively distilled from the teacher configuration and internalized within the student model’s parameters for inference.